### v1

# Safe Explicable Policy Search

Abstract
When users work with AI agents, they form conscious or subconscious expectations of them. Meeting user expectations is crucial for such agents to engage in successful interactions and teaming. However, users may form expectations of an agent that differ from the agent’s planned behaviors. These differences lead to the consideration of two separate decision models in the planning process to generate explicable behaviors. However, little has been done to incorporate safety considerations, especially in a learning setting. We present Safe Explicable Policy Search (SEPS), which aims to provide a learning approach to explicable behavior generation while minimizing the safety risk, both during and after learning. We formulate SEPS as a constrained optimization problem where the agent aims to maximize an explicability score subject to constraints on safety and a suboptimality criterion based on the agent’s model. SEPS innovatively combines the capabilities of Constrained Policy Optimization and Explicable Policy Search. We evaluate SEPS in safety-gym environments and with a physical robot experiment to show that it can learn explicable behaviors that adhere to the agent’s safety requirements and are efficient. Results show that SEPS can generate safe and explicable behaviors while ensuring a desired level of performance w.r.t. the agent’s objective, and has real-world relevance in human-AI teaming.
I Introduction
Recent developments in Artificial Intelligence (AI) have led to AI systems being widely deployed for user access. Users are increasingly involved in close interaction with these systems, forming a team-like relationship. In such cases, users form expectations of these “teammates”. For successful teaming interactions, it is important that user expectations are met. Otherwise, it may lead to user confusion, poor teaming performance, and loss of trust [1].
Generally, users may not always be experts in AI systems, which may lead to them forming expectations that are different from the system’s planned behaviors. These differences may be due to, for instance, different understandings of the task objective and/or domain dynamics during decision-making. Explicable planning [2, 3] addresses the planning challenges where the AI agent aims to generate more expected plans or policies under such model differences. However, when the agent focuses on meeting user expectations, it may lead to safety issues.
A previous study [4], introduced the problem of Safe Explicable Planning (SEP) that aims to find user expected policies that are safe by considering a safety bound (provided by an expert) in the agent’s decision making (Fig. 1(b)). This work is limited to discrete states and actions, and to a planning setting where the models of the agent and the user (for generating expectations of the agent) are known. In reality, it is desirable for such methods to work in environments with continuous states and actions, and, more importantly, when the user’s model is not directly available. Our work bridges the gap by proposing a method for finding safe and explicable policies in continuous stochastic environments in a reinforcement learning (RL) setting. It is worth noting that, in a planning setting, the solution is a final plan or policy that is safe. However, in a learning setting, the solution may require the agent to be safe even during the learning process.
(a) Motivating example
(b) Problem setting of SEP and the operation flow of SEPS
Figure 1: (a) The car merging scenario showing the user expected behavior (left), the optimal behavior (middle), and the safe explicable behavior (right), (b) The agent learns a surrogate reward function $u_{H}$ that captures the user’s expectations that are generated from her belief of the agent’s model $\mathcal{M}_{A}$, represented as $\mathcal{M}_{A}^{H}$. In a reinforcement learning setting, when the agent takes an action, it receives the agent’s task reward $\mathcal{R}_{A}$, the safety costs associated with a set of cost functions, $\mathcal{C}_{i}$, and the assumed surrogate reward $u_{H}$ as user feedback. The agent approximates the expected returns for these costs and rewards while applying the quality $d_{0}$ and safety limits $\{d_{i}\}$ to ensure that its policy is safe and efficient.
Consider a motivating example, shown in Fig. 1(a), where an autonomous driving car is to merge with highway traffic. The safety criterion is to maintain a minimum distance from both the car immediately in front and the car immediately behind. The AI driving agent aims at saving energy (by reducing acceleration and deceleration) while the human values travel time more. The expected behavior is to merge as fast as it can. Even when assuming safety is assured, it may result in sudden and harsh acceleration and/or deceleration when the front and/or rear cars are very close to the AI agent. The optimal behavior to the agent is to slow down to wait for a better opportunity (where more distance in between the AI agent and the other cars are observed) to smoothly accelerate and merge. The safe explicable behavior may involve alternating between deceleration and acceleration for an opportunity (with moderate distances between the AI agent and other cars) to merge that does not require too much waiting. The proposed Safe Explicable Policy Search (SEPS) aims to best match the user’s expectation within the agent’s safety and quality limits, thus avoiding unsafe and inefficient behaviors during and after learning.
Both safe behavior and explicable behavior generation, considered independently, can be addressed by the existing methods. For example, Constrained Markov Decision Process (CMDP) [5] and its solution techniques [6, 7, 8], etc., are effective in optimizing a task objective subject to safety constraints. The challenge in our problem is that the objective is defined over two different models with potentially different objectives and domain dynamics. This limits us from directly applying CMDP solutions. Similarly, a previous study on Explicable Policy Search (EPS) [3] proposes learning a surrogate reward function from user feedback on the agent’s behavior to inform learning. It is more practical than learning both the objective and domain dynamics and hence can be applied when the user’s model is unknown. However, EPS depends on the specification of a reconciliation parameter, and a weighting parameter for a few entropy terms, to combine an explicability score with the agent’s objective. It requires careful fine-tuning via trial and error due to the difficulty in relating these hyper-parameters to the importance of the objectives reconciled since they are at different scales (i.e., reward vs. entropy). We propose SEPS to combine the capabilities of CMDP (CPO [6]) and EPS while avoiding these limitations in addressing safe explicable behavior generation.
Towards this end, we first formulate our objective based on EPS since it allows us to convert the consideration of two models to the consideration of two linearly combined objectives. This is particularly beneficial since it enables us to focus policy search on a single model, thus allowing us to apply CPO to the new objective. Meanwhile, to relax the requirement of the hyper-parameters, we adapt and reformulate the objective by breaking it into an objective and a constraint, where the constraint can be absolved into the CPO formulation, leading to a novel integration of CPO and EPS while avoiding their limitations for safe explicable behavior generation. The remaining challenge lies in the derivation of an analytic solution for the resulting constrained optimization problem, for which we provide a solution for the most common case with two constraints. Furthermore, we evaluate SEPS in different safety-gym environments and with a physical robot experiment. The results show that our approach is effective in searching for an explicable policy within the safety limits. Our approach also demonstrates its applicability to real-life scenarios in the robot experiment.
II Related Work
With recent developments in AI, AI agents become accessible to various users where they either work alongside or interact with them. This has motivated advancements in Explainable AI (XAI) with the aim of building AI agents whose behavior is explainable to users [9, 10]. Our work is related to research that focuses on generating user-understandable behavior which is studied under different terminologies, viz., legibility [11], predictability [12], transparency [13], explicability [2], etc. For a detailed review of these studies, refer to [14].
In the spectrum of Explainable AI literature, our work is most related to explicable planning [2, 3, 15, 16], a framework for finding a user expected behavior while balancing the costs incurred to the agent. These studies propose different methods to measure similarity to user expectations but do not consider safety issues. Other studies that consider human feedback in the agent’s learning process, such as  [17, 18], are capable of learning user preferred behavior in the presence of an expert designed RL objective. However, these methods assume the user’s belief of the domain dynamics is the same as that of the agent, which can lead to learning undesired behaviors [19]. Furthermore, these approaches do not consider safety issues either.
Safe Explicable Planning (SEP) extends explicable planning to support the specification of a safety bound [4]. However, this work is limited to a planning setting where the decision models are given. In reality, such models may be difficult to provide a priori. Furthermore, it requires searching through the policy space, which is intractable in large, continuous MDPs. In this work, we consider the problem of finding a safe explicable policy in a reinforcement learning setting. However, we assume that the information regarding the user’s model is learned from user feedback and is given. This allows us to focus on developing the methodology for SEPS. Furthermore, unlike SEP, in our formulation, we consider the general case of continuous state and action spaces.
In our work, we define the agent’s decision model as a Constrained Markov Decision Process (CMDP)  [5], which allows us to specify the costs associated with safety separately from the agent’s task objective. CMDPs are popularly used in Safe RL (refer to [20] for a review of Safe RL methods) due to the ease of safety specification. Several techniques have been proposed to solve safe RL formulated as CMDPs [6, 21, 22, 23]. However, these methods cannot be directly applied to solve SEPS due to its problem setting (refer to Fig. 1(b)), where the optimization is under two different models, specifically different domain dynamics models (or transition models in MDPs).
III Preliminaries
A Markov Decision Process (MDP) is represented by a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho,\gamma\rangle$, where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the transition probability function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ is the reward function, $\rho(s)$ is the initial state distribution, and $\gamma$ is the discount factor. A stationary policy $\pi:\mathcal{S}\mapsto\mathcal{P}(\mathcal{A})$ is a map from states to probability distribution over actions, with $\pi(a|s)$ denoting the probability of selecting action $a$ in state $s$. We denote the set of all stationary policies by $\Pi$. The performance (expected discounted return) of a policy w.r.t. a measure $\mathcal{R}$ under the transition dynamics $\mathcal{T}$ is typically given by $J_{\mathcal{R},\mathcal{T}}(\pi)\dot{=}E_{\tau\sim\langle\pi,\mathcal{T}\rangle}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t},s_{t+1})]$, where $\tau\sim\langle\pi,\mathcal{T}\rangle$ indicates that the distribution over trajectories depends on $\pi$ and $\mathcal{T}$ i.e., $s_{0}\sim\rho,a_{t}\sim\pi(.|s_{t}),s_{t+1}\sim\mathcal{T}(.|s_{t},a_{t})$. The probability of realizing a trajectory $\tau$ with policy $\pi$ is $p(\tau){=}\rho(s_{0})\prod_{t}\mathcal{T}(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})$. Reinforcement learning algorithms (refer to [24]) seek to learn a policy that maximize the performance, i.e., $\pi^{*}=\operatorname{arg\,max}_{\pi\in\Pi}J_{\mathcal{R},\mathcal{T}}(\pi)$.
A constrained Markov Decision Process (CMDP) is an MDP augmented with constraints that restrict the permissible policies for the MDP. Specifically, $\mathcal{M}$ is augmented by auxiliary cost functions $\{\mathcal{C}_{i}\}_{i=1:k}$ and limits $\{d_{i}\}_{i=1:k}$ where each $\mathcal{C}_{i}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ is a function, same as $\mathcal{R}$. The performance of a policy w.r.t. $\mathcal{C}_{i}$ is $J_{\mathcal{C}_{i},\mathcal{T}}(\pi)=E_{\tau\sim\langle\pi,\mathcal{T}\rangle}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{C}_{i}(s_{t},a_{t},s_{t+1})]$. Then, the set of all feasible stationary policies is denoted by $\Pi_{\mathcal{C}}\dot{=}\{\pi{\in}\Pi:\forall{i}J_{\mathcal{C}_{i},\mathcal{T}}(\pi)\leq d_{i}$}. The objective of a CMDP is to learn a feasible policy that maximizes the performance $\pi^{*}=\operatorname{arg\,max}_{\pi\in\Pi_{\mathcal{C}}}J_{\mathcal{R},\mathcal{T}}(\pi)$.
In explicable policy search (EPS) [3], two MDP models, the agent’s model $M_{A}$ and the user’s belief of the agent’s model $M_{A}^{H}$, are considered at play. The two MDPs share the same $\mathcal{S}$, $\mathcal{A}$, $\rho$, and $\gamma$ but may differ in their domain dynamics ($\mathcal{T}_{A}$, $\mathcal{T}_{A}^{H}$) and their reward functions ($\mathcal{R}_{A}$, $\mathcal{R}_{A}^{H}$). In EPS, the aim is to search for a policy that maximizes two objectives: the expected cumulative reward, and a policy explicability score, linearly weighted by a reconciliation factor ($\lambda$) i.e.,
$\pi^{*}=\operatorname{arg\,max}_{\pi\in\Pi}J_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi)+\lambda\hskip 2.0pt\mathcal{E}(\pi,M_{A},\pi_{A}^{H},M_{A}^{H}).$ (1)
The explicability score (.$\mathcal{E}(.){=}-\mathcal{D}_{KL}(p_{A}(\tau)||p_{A}^{H}(\tau))$ is well motivated (refer to  [3]) and is defined as the negative KL-divergence of the trajectory distributions in the agent’s model and the human’s model denoted by $p_{A}(\tau)$ and $p_{A}^{H}(\tau)$, respectively. Essentially, the agent that maximizes $\mathcal{E}(.)$ would learn to replicate (as closely as possible) the human’s expected behavior of the agent.
Estimating the parameters in $p_{A}^{H}(\tau)$ (i.e. $\mathcal{T}_{A}^{H}$ and $\pi_{A}^{H}$), individually, is expensive. Instead, the authors propose to learn a surrogate reward function $u_{H}$ via a preference-based learning framework [25, 26] by presenting pairs of trajectories and asking human subjects which trajectory is expected. The learned reward function $u_{H}$ correlates with the distribution of human expected trajectories, $p_{A}^{H}(\tau)$. Furthermore, assuming a parametrized agent policy $\pi_{\theta}$, the surrogate reward signal and other log terms in the explicability measure effectively reshape the reward function in Eq. 1 and is approximated by the objective
$\theta^{*}\dot{=}\operatorname{arg\,max}_{\pi_{\theta}}\mathbb{E}_{p_{A}}\biggl{[}\sum_{t}\gamma^{t}(\mathcal{R}_{A}(s_{t},a_{t})\\ +\lambda(u_{H}(s_{t},a_{t})+\mathcal{H}_{\mathcal{T}_{A}}(s_{t+1}|s_{t},a_{t})+\mathcal{H}_{\pi_{\theta}}(a_{t}|s_{t})))\biggr{]},$ (2)
where $u_{H}(s_{t},a_{t})=\mathcal{T}_{A}^{H}(s_{t+1}|s_{t},a_{t})+\pi_{A}^{H}(a_{t}|s_{t})+C_{1}$, $\mathcal{H}_{\pi_{\theta}}$ is the policy entropy term (same as MaxEntRL [27]) which incentivizes the agent to explore uncertain actions, and $\mathcal{H}_{\mathcal{T}_{A}}$ is the environment entropy term which incentivizes the agent to explore parts of the environment that has more stochasticity.
IV Problem Formulation
In this section, we motivate and discuss the formulation of SEPS by incorporating the benefits of CMDP solution techniques (specifically, that of CPO [6]) with EPS [19]. Extending CPO to generate safe explicable policies is not straightforward due to the consideration of the objective over two different MDP models. We refer the reader to SEP [16], which discusses the challenges of solving a constrained optimization problem defined over two separate transition models. Similarly, extending EPS to generate safe explicable policies requires the specification of a reconciliation parameter, and a weighting parameter for a couple of entropy terms to scale the different measures with the reward appropriately. We propose a formulation for SEPS that systematically combines the capabilities of CPO and EPS while eliminating their limitations towards finding safe explicable policies.
Exhibiting explicable behavior is important for successful teaming. Examples of RL approaches that consider human feedback are [17, 18, 28, 3]. A popular objective for finding explicable plan or policy is to optimize an objective that linearly combines the agent’s rewards with an explicability measure [2, 3, 16]. However, being explicable at the cost of safety can be dangerous. Hence, it is unacceptable to compromise safety.
Ensuring safety in a task means avoiding or reducing harm. Implicitly, any task assumes successful completion without causing harm. For instance, if an agent is tasked with fetching hot coffee, an underlying safety assumption is that it does not bump into anyone on its way or spill the hot coffee. One approach, albeit nonideal, is to optimize an objective that balances risk penalties along with rewards, and explicability. This requires carefully scaling each component to provide a meaningful learning signal at every step, preventing issues like reward hacking. In safety-critical tasks, designing a single reward that simultaneously drives task completion and discourages unsafe behavior is challenging. A common approach is to define the reward as $\mathcal{R}=R-\omega C$, where $R$ represents the task reward component, and $C$ denotes the safety cost. However, results from CPO [6] demonstrate high sensitivity to the penalty coefficient $\omega$, highlighting the difficulty of crafting an effective overloaded reward function. Prior research [29], also emphasizes the importance of separating the task objective from safety constraints, as done in CMDPs, to enable the agent to learn meaningful behavior.
CMDP methods such as [5, 6, 7, 8], etc. efficiently optimize the task objective while enforcing safety constraints. Ideally, an agent should act both explicably (which involves considering two models) and efficiently while satisfying safety requirements. However, a CMDP requires both the objective and constraints to be defined under a single transition model. To address this, we propose to consider the objective in EPS, since it converts the consideration of two models to the consideration of two linearly combined objectives under a single transition model, as shown below,
$\pi^{*}=\operatorname{arg\,max}_{\pi\in\Pi_{d}}$ $J_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi)+\lambda\hskip 2.0pt\mathcal{E}(\pi,\mathcal{M}_{A},\pi_{A}^{H},\mathcal{M}_{A}^{H})$
$s.t.\hskip 3.0pt$ $J_{\mathcal{C}_{i},\mathcal{T}_{A}}(\pi)\leq d_{i}\hskip 10.0pti=1,....,k.$ (3)
This objective has two main limitations, 1) similar to the penalty coefficient, determining an appropriate reconciliation factor ($\lambda$) can be challenging, and 2) since the task objective is maximized alongside explicability, for $\lambda>0$, there is no guarantee of task performance in the agent’s model, especially when $\lambda$ is high. To overcome these issues, we propose to constrain the agent’s task objective instead of linearly combining it with the explicability measure as below.
$\pi^{*}=\operatorname{arg\,max}_{\pi\in\Pi_{d}}$ $\mathcal{E}(\pi,\mathcal{M}_{A},\pi_{A}^{H},\mathcal{M}_{A}^{H})$
$s.t.\hskip 3.0pt$ $J_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi)\geq d_{0}$
$J_{\mathcal{C}_{i},\mathcal{T}_{A}}(\pi)\leq d_{i}\hskip 10.0pti=1,....,k$ (4)
where $\mathcal{E}(.)$ can be written as $J_{\mathcal{E},\mathcal{T}_{A}}(\pi)$ based on Eq. 2. Essentially, $J_{\mathcal{E},\mathcal{T}_{A}}(\pi)\dot{=}E_{\tau\sim\langle\pi,\mathcal{T}_{A}\rangle}[\sum_{t=0}^{\infty}\gamma^{t}(u_{H}(s_{t},a_{t})+\mathcal{H}_{\mathcal{T}_{A}}(s_{t+1}|s_{t},a_{t})+\mathcal{H}_{\pi_{\theta}}(a_{t}|s_{t}))]$. Since the entropy terms ($\mathcal{H}_{\mathcal{T}_{A}}$ and $\mathcal{H}_{\pi_{\theta}}$) are approximated separately from the surrogate reward ($u_{H}$), they require additional scaling parameters which are not straightforward to determine. In our work, we mitigate this by simply maximizing $u_{H}$. This means that instead of maximizing the negative KL divergence between the trajectory distributions, we maximize the probability of the most expected trajectory that the agent can generate.
First, let us define the components involved in formulating SEPS. The agent’s decision model is modeled as a CMDP augmented with an additional limit on the task objective i.e., $\mathcal{M}_{A}=\langle\mathcal{S},\mathcal{A},\mathcal{T}_{A},\mathcal{R}_{A},\{\mathcal{C}_{i}\}_{i=1:k},\{d_{i}\}_{i=0:k},\rho,\gamma\rangle$ where, $d_{0}$ is the limit on the task objective $\mathcal{R}_{A}$ and $d_{1},....,d_{k}$ are limits on the cost functions $\mathcal{C}_{1},....,\mathcal{C}_{k}$, respectively. The user’s belief of the agent’s model is modeled as an MDP i.e., $\mathcal{M}_{A}^{H}=\langle\mathcal{S},\mathcal{A},\mathcal{T}_{A}^{H},\mathcal{R}_{A}^{H},\rho,\gamma\rangle$. Given the two decision models, we define SEPS as below.
Definition 1
Safe Explicable Policy Search (SEPS) is the problem of searching for a policy that maximizes the probability of the user’s expected trajectory subject to the constraints on the task objective and the costs associated with safety in the agent’s model, or formally,
$\pi^{*}\dot{=}\operatorname{arg\,max}_{\pi\in\Pi_{d}}$ $J_{u_{H},\mathcal{T}_{A}}(\pi)$
$s.t.\hskip 3.0ptJ_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi)$ $\geq d_{0}$
$J_{\mathcal{C}_{i},\mathcal{T}_{A}}(\pi)$ $\leq d_{i}\hskip 10.0pti=1,...,k.$ (5)
It is interesting to note that in the special case where $\mathcal{T}_{A}$ and $\mathcal{T}_{A}^{H}$ are identical, $u_{H}$ closely approximates the user’s reward model $\mathcal{R}_{A}^{H}$. However, we note that the goal of learning $u_{H}$ is to obtain feedback on the user’s assessment of explicableness (based on his own model - $\mathcal{T}_{A}^{H}$, $\mathcal{R}_{A}^{H}$) of agent’s behavior, i.e., when the agent operates under the true domain dynamics $\mathcal{T}_{A}$. Therefore, the SEPS problem in Eq. 5 is equivalent to the Safe Explicable Planning problem [4]
$\pi^{*}=\operatorname{arg\,max}_{\pi\in\Pi_{d}}$ $J_{\mathcal{R}_{A}^{H},\mathcal{T}_{A}^{H}}(\pi)$
$s.t.\hskip 3.0ptJ_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi)$ $\geq d_{0}$
$J_{\mathcal{C}_{i},\mathcal{T}_{A}}(\pi)$ $\leq d_{i}\hskip 10.0pti=1,...,k.$ (6)
V Safe Explicable Policy Search
A previous study on Safe Explicable Planning (SEP) [4] proposes searching through the policy space while evaluating policies in the agent’s model for feasibility and in the user’s model for explicablity. The SEP algorithms iterate over all states and actions to find an optimal, safe explicable policy, which is impractical in continuous MDPs. The SEPS problem in Eq. 5 reduces to a CMDP, allowing us to use its policy search techniques. CPO [6] is a popular approach to solving CMDP as it ensures safe behavior even during the learning process. However, the algorithm is limited to solving an objective with a single constraint. Therefore, we require an analytic solution to efficiently solve SEPS. As the number of constraints increases, deriving feasible bounds for the solution becomes challenging due to the intersection of feasible regions of every constraint.
Let us introduce some useful notations. The on-policy value function and action-value function in the agent’s model, over a general reward function $\mathcal{R}$, are given by $V_{\mathcal{R}}^{\pi}(s)\dot{=}E_{\tau\sim\langle\pi,\mathcal{T}_{A}\rangle}[\sum_{t}\gamma^{t}\mathcal{R}(s_{t},a_{t})|s_{0}{=}s]$ and $Q_{\mathcal{R}}^{\pi}(s,a)\dot{=}E_{\tau\sim\langle\pi,\mathcal{T}_{A}\rangle}[\sum_{t}\gamma^{t}\mathcal{R}(s_{t},a_{t})|s_{0}{=}s,a_{0}{=}a]$, respectively. Then, the corresponding advantage function is $A_{\mathcal{R}}^{\pi}(s,a)\dot{=}Q_{\mathcal{R}}^{\pi}(s,a)-V_{\mathcal{R}}^{\pi}(s)$. The function $\mathcal{R}$ may be replaced by $u_{H}$, $\mathcal{R}_{A}$, or any $\mathcal{C}_{i}$.
CPO applies trust region policy optimization (TRPO) [29] to policy search updates in a CMDP. Using the notation above, the formulation in Eq. 5 can be written as a policy update rule given by
[…]
where, $\Pi_{\theta}\subset\Pi$ is a set of parameterized policies with parameters $\theta$, $\bar{\mathcal{D}}_{KL}(\pi||\pi_{n})=\mathbb{E}_{s\sim\pi^{n}}[\mathcal{D}_{KL}(\pi||\pi_{n})[s]]$ is a measure of KL Divergence between the current policy $\pi_{n}$ and a potential new policy $\pi$, and $\delta>0$ is the step size. The set of policies that satisfy the third constraint is called the trust region, i.e. $\{\pi_{\theta}\in\Pi_{\theta}|\bar{\mathcal{D}}_{KL}(\pi||\pi_{n})\leq\delta\}$. For performance guarantees and constraint satisfaction guarantees, see [29] and [6], respectively.
(a) Returns under $u_{H}$
(b) Returns under $\mathcal{R}_{A}$
(c) Returns under $\mathcal{C}_{1}$
Figure 2: Average performance over three runs for SEPS and baselines in PointGoal; the x-axis is training epoch.
For continuous MDPs with high dimensional spaces, solving Eq. 7 exactly is intractable and requires an approximation. The CPO method proposes the following approximation
$\theta_{n+1}=\operatorname{arg\,max}_{\theta}g^{T}(\theta-\theta_{n})\\ s.t.\hskip 3.0ptc_{0}-b_{0}^{T}(\theta-\theta_{n})\leq 0\hskip 70.0pt\\ c_{i}+b_{i}^{T}(\theta-\theta_{n})\leq 0\hskip 10.0pti=1,...,k\\ \frac{1}{2}(\theta-\theta_{n})^{T}H(\theta-\theta_{n})\leq\delta\hskip 80.0pt$ (8)
where, $g$ is the gradient of the objective (explicability measure), $b_{0}$ is the gradient of the constraint on agent’s task objective, $b_{i}$ is the gradient of the constraint $i$, $c_{0}\dot{=}-J_{\mathcal{R}_{A},\mathcal{T}_{A}}(\pi_{n})+d_{0}$, $c_{i}\dot{=}J_{\mathcal{C}_{i},\mathcal{T}_{A}}(\pi_{n})-d_{i}$, and $H$ is the Hessian of the KL Divergence. The optimization problem mentioned above is convex. For policies with high-dimensional parameter spaces like neural networks, the primal problem above is impractical to solve (even with few constraints) using any convex program solver due to the computation complexity of the Hessian matrix.
Feasible case: When the problem is feasible the dual to the above is easier to solve and is given by
$\max_{\lambda\geq 0,\nu\succeq 0}\frac{-1}{2\lambda}(g^{T}H^{-1}g+2r^{T}\nu+\nu^{T}S\nu)+\nu^{T}c-\lambda\delta$ (9)
where $c\dot{=}[c_{0},c_{1},...,c_{k}]^{T}$, $B\dot{=}[b_{0},b_{1},...,b_{k}]$, $r\dot{=}g^{T}H^{-1}B$, and $S\dot{=}B^{T}H^{-1}B$. Here, the product terms involving the inverse of the Hessian matrix can be approximately computed as in [29]. If $\lambda^{*}$ and $\nu^{*}$ are optimal solutions to the dual above. The update rule for the feasible case is
$\theta_{n+1}=\theta_{n}-\frac{1}{\lambda}H^{-1}\left(g+\nu B\right)$ (10)
Infeasible case: When the policy $\pi_{\theta_{n}}$ is infeasible, due to a bad step taken previously (occurs due to approximation), the system can recover by updating the policy parameters to strictly reduce the constraint violation. In this case, the update rule is found by solving the following optimization problem depending upon the constraint being violated
$\theta_{n+1}=\operatorname{arg\,min}_{\theta}\hskip 5.0ptc_{m}-b_{m}^{T}(\theta-\theta_{n})\hskip 15.0pt\text{where, }m\in[0,k]\\ s.t.\,\,\,\,\frac{1}{2}(\theta-\theta_{n})^{T}H(\theta-\theta_{n})\leq\delta\hskip 50.0pt$ (11)
The dual to the above is
$\max_{\phi\geq 0}-\frac{s}{2\phi}+c_{m}-\phi\delta.$ (12)
where $s=b_{m}^{T}H^{-1}b_{m}$. If $\phi^{*}$ is the solution to the dual above. The update rule for the infeasible case is
$\theta_{n+1}=\theta_{n}-\frac{1}{\phi}H^{-1}b_{m}$ (13)
When multiple constraints are violated, we solve Eq. 11 for each violating constraint and combine the solutions in Eq. 13. This adjusts the policy towards the feasible region w.r.t. all constraints in a single step. Alternatively, one could update the policy with respect to only one constraint at a time.
We derive an analytical solution to SEPS for both feasible and infeasible cases above (refer to Theorems 1 and 2 in the extended version of this paper) when a single cost function is present, resulting in two constraints.
VI Evaluation
We evaluated our method in continuous safety-gym environments [30] and with a physical robot experiment. We designed the evaluation scenarios, including the risk elements. In all of the evaluations, we assume a learned surrogate reward function that captures SEPS relevant information regarding the user model. The goal of our evaluation is 1) to validate the efficacy of our method to search for safe explicable policies which a) meet user expectations, and b) are safe and efficient, 2) to analyze the behaviors generated by our method in comparison with that in the agent’s model and user expectations, and 3) to demonstrate the applicability of SEPS to real-world scenarios.
VI-A Safety
In our first evaluation, we aim to answer ‘What are the benefits of safety constraints in explicable planning?’ and ‘Is SEPS’s policy explicable under safety constraints?’. We design an evaluation scenario such that the user’s expectations conflict with the safety requirements of the agent.
(a) Returns under $u_{H}$
(b) Returns under $\mathcal{R}_{A}$
(c) Returns under $\mathcal{C}_{1}$
Figure 3: Average performance over three runs for SEPS and baselines in PointButton task over random seeds; the x-axis is training epoch.
We conduct this evaluation on the safety-gym’s PointGoal task as shown in Fig. 4(a). Here, the agent is required to reach the goal (green) while avoiding stepping into hazardous areas (dark blue) and hitting vases (light blue). In the agent’s model, the task objective is to reach the goal efficiently. Furthermore, there are constraints on the costs associated with hazards and vases. We assume the user’s belief of the agent’s objective includes avoiding moving close to vases as they are considered fragile. However, the user is unaware of the safety risk associated with the agent entering the hazardous areas.
We evaluated SEPS against the agent’s performance in its own model ($\mathcal{M}_{A}$), under user expectations ($u_{H}$), and under EPS ($EPS$). Here, no safety constraints are incorporated into the baselines $u_{H}$ and $EPS$, to emphasize the need for safety constraints in explicable planning. In the next evaluation, we will incorporate the consideration of safety constraint into these baselines to compare with SEPS. We compared the behaviors generated and the returns obtained under the surrogate reward, the agent’s task reward, and the cost function.
Results and Discussion
The results of this evaluation are shown in Fig. 2. The main result is in the violation of safety constraints reflected in the returns under the cost function $\mathcal{C}_{1}$ shown in Fig. 2(c). The closer the cost return to 00, the safer the behavior is. $u_{H}$ and EPS resulted in more safety violations than SEPS and the optimal behavior (which happened to incur a low cost due to the environment setting). This is due to the fact that the surrogate objective only includes consideration of the vases. EPS incurred the most cost since it sometimes resulted in the agent wandering around without reaching the goal, leading to the low return in Fig. 3(b) and demonstrating the instability issue due to linearly combining objectives. Even though the agent in SEPS executed a curvy path to the goal, its return in $\mathcal{R}_{A}$ is almost as high as the optimal behavior (Fig. 3(b)). This is because the task objective in the agent’s model features a penalty proportional to its displacement from the goal. As long as the agent makes progress towards the goal (whether straightly or in a curve), little penalty would be given.
(a) PointGoal
(b) PointButton
Figure 4: Lines show resulting trajectories of different methods in the two simulation domains; SEPS-blue, $\mathcal{M}_{A}$/$\mathcal{R}_{A}{-}CPO$-red, $u_{H}$/$u_{H}{-}CPO$-orange, and $EPS$/$EPS{-}CPO$-green.
Fig. 4(a) shows the behaviors generated in the PointGoal task. As user model is assumed to not associate risk with hazardous regions but only with vases, $u_{H}$ and $EPS$ generated unsafe trajectories that passed through hazardous regions. In $\mathcal{M}_{A}$, the agent followed the most efficient path to reach the goal without colliding with the vases or entering the hazardous regions, although such a path narrowly passing through the two vases may be seen as risky. $SEPS$ also avoided hazardous regions and staying further away from the vases, to be explicable.
VI-B Efficiency
In our second evaluation, our aim is to answer ‘Is the agent efficient under its own task objective?’ and ‘Is SEPS’s policy explicable and safe while being efficient?’. We designed an evaluation scenario where the user’s model aligned with the safety requirements of the agent but differed from the agent’s task objective.
We conducted this evaluation in a modified version of the safety-gym’s PointButton task as shown in Fig. 4(b). Here, the agent is required to reach the goal (green) while avoiding contact with moving gremlins (purple). In addition, there are two buttons (orange) in the environment that can be pressed for a small reward. In the agent’s model, the task objective is to reach the goal efficiently subject to constraints on the costs associated with gremlins. We assumed that the user’s belief of the agent’s objective was to press all the buttons to maximize its reward before reaching the goal while avoiding the gremlins.
We evaluated SEPS against three baselines that enforce safety constraints via CPO: the agent’s objective $\mathcal{R}_{A}-CPO$, the user’s expectations (surrogate reward) $u_{H}-CPO$, and Explicable Policy Search $EPS-CPO$. We compared the behaviors generated and the returns obtained under the surrogate reward, the agent’s task reward, and the cost function.
Results and Discussion
The results for this evaluation are shown in Fig. 3. As $u_{H}-CPO$ solely optimizes the surrogate objective ($u_{H}$), it performed poorly under the agent’s task objective ($\mathcal{R}_{A}$). Similarly, as $\mathcal{R}_{A}-CPO$ solely optimizes $\mathcal{R}_{A}$, it performed poorly under $u_{H}$. Explicable Policy Search (EPS), when optimized subject to safety constraints ($EPS-CPO$), achieved better performance under the surrogate reward but performed poorly under the agent’s task objective due to the lack of any constraint on the objective. In $EPS-CPO$ we set the reconciliation parameter $\lambda=3$. Empericially, we found that $\lambda=3$ balanced reasonably well between the surrogate reward and the agent’s reward.
(a) $\tau_{\mathcal{M}_{A}}$
(b) $\tau_{\mathcal{M}_{A}^{H}}$
Figure 5: Robot optimal behavior and explicable behavior.
In $SEPS$, we set $d_{0}=0.0$ and $d_{1}=2.5$. Note that these bounds have clear semantic meanings (e.g., the suboptimality level under the respective metric), which makes them easier to specify than the hyper-parameters in EPS. It is worth noting that unlike $SEPS$, all baselines have a single constraint on safety. Addressing multiple constraints is more challenging since they may sometimes conflict with each other. Even under such a case, $SEPS$ stayed close to $\mathcal{R}_{A}-CPO$ under both $\mathcal{R}_{A}$ and the safety cost. In other words, it is able to remain explicable while sufficiently satisfying both the performance and safety requirements.
The behaviors generated by different methods for an example seed in the PointButton task are shown in Fig. 4(b). In SEPS and $EPS-CPO$ the agent pressed a single button before reaching the goal. However, in $EPS-CPO$ the agent showed a suboptimal behavior by taking a longer path to the button and going around the button before reaching the goal. As expected, in $u_{H}-CPO$ the agent pressed both buttons while in $\mathcal{R}_{A}-CPO$ it headed straight to the goal.
VI-C Physical Robot Experiment
In this experiment, our aim is to analyze the behaviors generated when a physical robot (UR5) assists a user to set up a dining table. We designed the experiment (see Fig. 5) in a discrete setting where the location of the objects was preset to specific waypoints, and the robot’s movement between them was calibrated using the UR5’s inverse kinematics library.
We assume that the user’s expectations differ from the robot’s functionality. The robot is required to pass the cups to the user by placing them on the dining table. The user expects the robot to place the cups close to where he is standing. However, doing so may tip the glass in front of the robot. Tipping the glass will spill its contents, which is unsafe in the robot’s model and has a high cost. We analyzed the behavior generated by SEPS and compared it with the optimal behavior in its own model, the user’s expected behavior, and $EPS{-}CPO$.
(a) $\tau_{EPS}$
(b) $\tau_{SEPS}$
Figure 6: Safe explicable behaviors generated by $EPS{-}CPO$ and SEPS.
Results and Discussion
The robot picked the cup and placed it on the table next to the glass when executing the optimal behavior in its model (see Fig. 5(a)). When executing the user’s expected behavior, the robot tipped the glass (replaced by an empty plastic bottle in this execution for safety reasons), as shown in Fig. 5(b). In reality, the robot would have spilled the liquid as the glass obstructed the manipulator. With EPS (see Fig. 6(a)), the robot first picked up the glass and moved it across the table. Then, it picked up the cup and placed it close to the user to meet his expectation. However, picking and placing the glass was less efficient due to additional movements. With SEPS (see Fig. 6(b)), the robot first moved the manipulator to push the glass gently across the table to clear it out of the way. Then, it picked up the cup and placed it close to the user as in EPS. While both EPS and SEPS generated explicable behaviors, SEPS’s is more efficient, thanks to the constraint on the task objective.
VII Conclusion
In this paper, we introduced the problem of Safe Explicable Policy Search (SEPS), which aimed to search for explicable policies while considering both safety and task efficiency. We adopted and adapted the explicability measure in EPS to formulate SEPS as a CMDP. SEPS relaxed the need for specifying multiple hyper-parameters for explicable policy search by considering the agent’s task objective as a separate constraint. These treatments allowed us to apply CPO to solve SEPS to satisfy the safety and quality requirements during the learning process while maximizing explicability. We derived an analytical solution for the most common case of SEPS with two constraints. We evaluated SEPS for safety and efficiency in safety-gym environments. The results demonstrated that SEPS was effective in learning safe, explicable, and efficient policies. We also demonstrated the applicability of our approach via a physical robot experiment.